OpenStudy (anonymous):

to make an internacional telephone call. you need the code for the country you are calling. The codes for country A, coun try B, and C are three consecutive integers whose sum is 93. Find the code for each country Country's A code is , country's B code is , and country;s C code is .

OpenStudy (stormfire1):

The equation would look like: \[n+(n+1)+(n+2)=93\] Which reduces to 3n=90, so n = 30, n+1 = 31, n+2 = 32.